# DERD-Net: Learning Depth from Event-based Ray Densities

### v1

@@ Appendix A Supplementary Material
The averaged results of different SOTA methods compared to DERD-Net are presented in Tab. 8. In Tabs. 9, 10 and 11, the individual performance on each of the respective sequences indoor_flying 1, 2, 3 from the MVSEC dataset are displayed. Table 12 presents the corresponding results for the Zurich_City_04_a sequence from the DSEC dataset. For the data augmentation and sensitivity analysis, the full metrics are disclosed in Tabs. 13 and 14, respectively.
@@
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo ESVO indep. 1s stereo SGM indep. 1s stereo GTS indep. 1s stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
@@
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [m] [m] [%] [%] [%] [%] [%] [million] SOTA EMVS [31] monocular + EMVS [31] monocular + ESVO [46] stereo MC-EMVS [13] stereo + MC-EMVS [13] stereo + MC-EMVS [13] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
Data Augmentation Analysis
Sensitivity Analysis

### v2

@@ Appendix B Appendix - Detailed per-Sequence Results
The average results of different SOTA methods compared to DERD-Net are presented in Tab.˜11. In Tabs.˜12, 13 and 14, the individual performance on each of the respective sequences indoor_flying 1, 2, 3 from the MVSEC dataset are displayed. Table˜15 presents the corresponding results for the Zurich_City_04_a sequence from the DSEC dataset.
@@
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [cm] [cm] [%] [%] [%] [%] [%] [million] SOTA EMVS [11] monocular + EMVS [11] monocular + ESVO [19] stereo ESVO indep. 1s stereo SGM indep. 1s stereo GTS indep. 1s stereo MC-EMVS [12] stereo + MC-EMVS [12] stereo + MC-EMVS [12] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net without EL stereo + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +
@@
Algorithm Modality Mean Err Median Err bad-pix SILog Err AErrR log RMSE #Points [m] [m] [%] [%] [%] [%] [%] [million] SOTA EMVS [11] monocular + EMVS [11] monocular + ESVO [19] stereo SGM [38] stereo GTS [39] stereo MC-EMVS [12] stereo + MC-EMVS [12] stereo + MC-EMVS [12] + MF stereo + Ours DERD-Net monocular + DERD-Net monocular + DERD-Net stereo + DERD-Net stereo + DERD-Net multi-pixel stereo + DERD-Net multi-pixel stereo +